OpenStudy (anonymous):

Joe chops a 12 foot long log into 3/4 pieces. How many sections of firewood does he end up with?

OpenStudy (anonymous):

@ghazi

OpenStudy (anonymous):

@agent0smith

OpenStudy (agent0smith):

Uh, is it 3/4 foot pieces? 3/4 inch? Can't answer it without knowing that.

OpenStudy (anonymous):

I did 12/3/4 on wrb20calc and got 1

OpenStudy (anonymous):

oh 3/4 foot

OpenStudy (agent0smith):

So you're going at it the right way, 12/(3/4) but you've made a mistake.\[\frac{ 12 }{ \frac{ 3 }{ 4} }=12 \times \frac{ 4 }{ 3 }\]

OpenStudy (agent0smith):

When you divide by a fraction, you can multiply by the reciprocal of the fraction.

OpenStudy (anonymous):

oh. so. is it 1?

OpenStudy (agent0smith):

it won't be 1, think about it this way - if you have a 12 foot long piece of wood, and you cut it into 1 foot long pieces, how many pieces would you have? 12/1 right? So 12. And 3/4 foot is smaller than 1 foot, so there will be more than 12 pieces.

OpenStudy (anonymous):

.oh.

OpenStudy (anonymous):

so how would u work it out?

OpenStudy (agent0smith):

\[\frac{ 12 }{ \frac{ 3 }{ 4} }=12 \times \frac{ 4 }{ 3 } =\frac{ 12 \times 4 }{ 3 }\]

OpenStudy (anonymous):

so do i calculate that?

OpenStudy (agent0smith):

yep

OpenStudy (anonymous):

mk. brb

OpenStudy (anonymous):

16?

OpenStudy (agent0smith):

Correct. You can do it on a calculator, but the problem is that you forgot the brackets. "I did 12/3/4 on wrb20calc and got 1" 12/3/4 is NOT the same as 12/(3/4) <---- enter this on a calculator and you'll get the correct answer. Enter 12/3/4 and you'll get 1.

OpenStudy (anonymous):

ah. so, 16 ir right,yes?

OpenStudy (agent0smith):

Yes. You can check it on a calculator, by entering 12/(3/4), just don't forget the brackets.

OpenStudy (anonymous):

i did that just now. tnx:)

OpenStudy (agent0smith):

Or, convert 3/4 to a decimal = 0.75. Then do 12/0.75 = 16.

OpenStudy (anonymous):

okay.